### v1

# Measurement of the longitudinal flow-plane decorrelation using multi-plane cumulants in $\sqrt{s_{{}_{\mathrm{NN}}}}$ = 200 GeV Au+Au, Ru+Ru, and Zr+Zr collisions

Abstract
Measurements of the variation of anisotropic flow-plane angles ($\Psi_{n}$) with rapidity, commonly known as the flow-plane decorrelation, provide important insights into the initial conditions of the matter produced in heavy-ion collisions. In this paper, using data collected by the STAR experiment, we report the first measurement of the four-plane cumulant observable $T_{n}\{ba;dc\}=\langle\langle\sin[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle\rangle$, where superscripts $a$, $b$, $c$, and $d$ denote sequential regions in pseudorapidity with $a$ corresponding to the most backward pseudorapidity region. The measurement is performed for the elliptic and triangular flow (i.e. $n=2$ and $3$) in Au+Au and isobar (Ru+Ru, Zr+Zr) collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}$ = 200 GeV. The goal of calculating the correlation of the flow-plane angle variations, from backward pseudorapidity ($\eta_{a}$) to mid-pseudorapidity ($\eta_{b}$) region and from mid-pseudorapidity $(\eta_{c}$) to forward pseudorapidity ($\eta_{d}$) region, is to probe the systematic variation of flow angle over a wide pseudorapidity range. In mid-central collisions ($10-30\%$ centrality), we find $T_{2}\{ba;dc\}=-0.004\pm 0.001(\rm stat)\pm 0.002(\rm syst)$ independent of the collision system. Such a small value of $T_{2}$ favors a “random-walk” variation of the flow-plane angles, where the rapidity correlation length is smaller than the pseudorapidity region under study. These measurements provide new information on the decorrelation patterns in the system and offer a quantitative estimate of possible systematic variations in anisotropic flow angles-such as “twist” between forward and backward pseudorapidity regions. This opens new opportunities for understanding the three-dimensional structure of the quark-gluon plasma created in heavy-ion collisions.
pacs:
25.75.-q, 25.75.Ld, 25.75.Gz
I Introduction
The STAR experiment at the Relativistic Heavy Ion Collider (RHIC) at Brookhaven National Laboratory (BNL) focuses on investigating the formation of a deconfined state of quarks and gluons known as the quark-gluon plasma (QGP) et al. (STAR Collaboration); Busza et al. (2018). One of the primary goals of STAR and other heavy ion experiments is to probe the mechanism behind the QGP formation and to understand its evolution and hadronization processes. It has long been realized that the QGP created in such collisions exhibits significant initial-state spatial anisotropies which originate from both the geometry of non-central nuclear collisions and the probabilistic nature of nucleon-nucleon interactions. The spatial anisotropies in the initial-state matter distribution lead to momentum anisotropies of the final-state particles due to the space-momentum interactions during system evolution.
The azimuthal anisotropy in the distribution of emitted particles can be characterized by a superposition of different modes or harmonics of collective flow in a Fourier series Voloshin and Zhang (1996); Poskanzer and Voloshin (1998); Voloshin et al. (2010)
$\frac{dN}{d\phi}\propto 1+2\sum_{n}^{\infty}v_{n}\cos[n(\phi-\Psi_{n})],$ (1)
where $\phi$ is the azimuthal emission angle of the particle, $v_{n}$ characterizes the strength of the $n$-th harmonic flow, and $\Psi_{n}$ is the flow (symmetry) plane angle. Flow coefficients $v_{1}$, $v_{2}$, and $v_{3}$ are often referred to as “directed”, “elliptic”, and “triangular” flow, respectively.
The reaction plane angle $\Psi_{\rm RP}$ is defined by the direction of the impact parameter vector. Using $\Psi_{\rm RP}$ in place of $\Psi_{n}$ in the decomposition Eq. 1 for all harmonics, corresponds to the flow ($v_{n}$) definition in the reaction-plane anisotropic flow picture. While this picture is well defined theoretically, in practice it is not very useful, as the reaction-plane angle can not be measured experimentally. In most cases, the flow itself (e.g., anisotropic particle distribution) is used to determine the flow-symmetry planes $\Psi_{n}$ as those corresponding to the maximum amplitude of a given harmonic. Due to the initial-state geometry fluctuations, flow planes $\Psi_{n}$ in general do not coincide with the reaction plane Alver and Roland (2010). In this so-called participant picture of flow, the flow development is determined by the initial geometry of the system, including fluctuations caused by the probabilistic character of the interactions of nucleons inside the colliding nuclei. The participant-plane description is meaningful only if the flow angles depend relatively weakly on the particle rapidity and, more generally, on the particle transverse momentum ($p_{\mathrm{T}}$) as well as particle type. Otherwise, the Fourier decomposition in Eq. 1 becomes impractical.
The study of the flow-plane angle $\Psi_{n}$, along with its fluctuations and correlations across rapidity, is of particular interest. Equally important are similar measurements of the flow coefficients $v_{n}$. These quantities are determined by fluctuations in the initial matter distribution. Consequently, they provide insights into various aspects of nuclear interactions, such as baryon stopping, shadowing, and sub-nucleon structure. Such fluctuations or variations of the flow plane with rapidity are referred to as longitudinal flow-plane decorrelations.
For discussion purposes, it is useful to separate the different sources of flow-plane decorrelations in rapidity. Short-range effects, such as jet energy deposition, influence only a limited region in rapidity. Long-range effects, on the other hand, arise from correlations like those in the “twisted fireball” scenario Bożek et al. (2011); Behera et al. (2020), where the flow plane in the forward region is primarily determined by the participant distribution from the projectile nucleus, while at backward rapidity, it is governed by the participants of target nucleus. The short-range effects when measured on a distance scale much larger than the correlation length appear as a “random-walk” variations in the flow angle. The resulting decorrelation is referred to as a random tilt. The “twisted fireball” scenario, on the other hand, can be visualized as a monotonic change in the flow-plane angle with rapidity. Another scenario corresponding to the long-range variation of flow plane occurs when the flow-plane angles in the forward and backward regions are close to each other, while those at the central region experiences relatively large fluctuations. This case would correspond to a non-monotonic change in the flow-plane angle with rapidity. See Refs. Pang et al. (2015, 2016); Wu et al. (2019); Jia and Huo (2014a, b); Xiao et al. (2013); Jia et al. (2024) for a more detailed discussion on the mechanisms of the flow-plane decorrelations.
Over the past decade, to study the flow decorrelation, different collaborations et al. (2015, 2018, 2021) often used an observable known as the factorization ratio:
$\begin{split}r_{n}(\eta_{a\rightarrow c/b})&\equiv\frac{\langle\cos[n(\phi^{c}-\Psi_{n}^{a})]\rangle}{\langle\cos[n(\phi^{b}-\Psi_{n}^{a})]\rangle}\\ &=\frac{\langle v_{n}(\eta_{c})\cos[n(\Psi_{n}^{c}-\Psi_{n}^{a})]\rangle}{\langle v_{n}(\eta_{b})\cos[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\rangle}\\ &\approx\frac{\langle\cos[n(\Psi_{n}^{c}-\Psi_{n}^{a})]\rangle}{\langle\cos[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\rangle}\\ &\equiv r^{\Psi}_{n}(\eta_{a\rightarrow c/b}),\end{split}$ (2)
where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the average over events. $\phi^{b}$ and $\phi^{c}$ are the azimuthal angles of particles from the two symmetric mid-pseudorapidity windows $b$ and $c$, respectively, on the backward and forward regions (see Fig. 1). Similarly, $\Psi_{n}^{a}$, $\Psi_{n}^{b}$, and $\Psi_{n}^{c}$ are the flow-plane angles of $n^{\rm th}$-order harmonic flow obtained from the backward pseudorapidity reference window $a$ ($\eta_{a}<0$) and two mid-pseudorapidity windows $b$ ($\eta_{b}<0$) and $c$ ($\eta_{c}>0$), respectively. The notations $\eta_{a}$, $\eta_{b}$, and $\eta_{c}$ refer to the ranges in pseudorapidity of three windows such that the reference window $a$ in the backward region is sufficiently far from the two mid-pseudorapidity windows $b$ and $c$ to introduce large pseudorapidity gaps, $|\eta_{b}-\eta_{a}|$ and $|\eta_{c}-\eta_{a}|$, in the measurement. The reference window in Eq. 2 can also be selected from the forward region (such as window $d$ in Fig. 1), provided that the mid-pseudorapidity regions $b$ and $c$ are simultaneously interchanged in the definition.
In symmetric collisions, such as Au+Au or Ru+Ru (Zr+Zr), the flow coefficients $v_{n}(\eta_{b})$ and $v_{n}(\eta_{c})$, corresponding to symmetric negative and positive pseudorapidity ranges $\eta_{b}$ and $\eta_{c}$, respectively, are expected to be identical when averaged over all events. Hence, the deviation of the ratio in Eq. 2 from unity would indicate the presence of flow decorrelation in the collisions. The observable $r_{n}(\eta_{a\rightarrow c/b})$ in Eq. 2 includes contributions from the decorrelations in the flow angles ($r_{n}^{\Psi}(\eta_{a\rightarrow c/b})$) as well as in the flow magnitudes, which is the reason for an approximate sign in the third line. As shown in Ref. Bozek and Broniowski (2018), the decorrelations in the flow magnitudes are relatively small, at the level of a few percent. However, the effect of using the flow magnitude weight in the calculation of the flow-plane decorrelation can be rather significant and comparable to the flow-angle decorrelation itself. This is due to the fact that the flow-plane fluctuations are smaller in the events with larger flow magnitude.
The observable $r_{n}(\eta_{a\rightarrow c/b})$ is sensitive to both the random tilt and the long-range systematic rotation of flow planes as in the “twisted fireball” scenario. As shown in Ref. Xu et al. (2022), the ratio $r_{n}^{\Psi}(\eta_{a\rightarrow c/b})$ is independent of the reference rapidity window $a$ in the random-walk scenario of the flow-plane angle variations.
In the small angle approximation, the observable $r_{n}^{\Psi}(\eta_{a\rightarrow c/b})$ can be expressed as
$\begin{split}&r^{\Psi}_{n}(\eta_{a\rightarrow c/b})\\ &\approx 1-\frac{n^{2}}{2}\langle(\Delta\Psi_{n}^{b\rightarrow c})^{2}\rangle-n^{2}\langle\Delta\Psi_{n}^{a\rightarrow b}\cdot\Delta\Psi_{n}^{b\rightarrow c}\rangle,\end{split}$ (3)
using the relationships
$\begin{split}\Delta\Psi_{n}^{a\rightarrow c}&=\Delta\Psi_{n}^{a\rightarrow b}+\Delta\Psi_{n}^{b\rightarrow c},\\ \Delta\Psi_{n}^{a\rightarrow c}&=\Psi_{n}^{c}-\Psi_{n}^{a},\\ \Delta\Psi_{n}^{a\rightarrow b}&=\Psi_{n}^{b}-\Psi_{n}^{a},\\ \Delta\Psi_{n}^{b\rightarrow c}&=\Psi_{n}^{c}-\Psi_{n}^{b},\end{split}$ (4)
where $\Delta\Psi_{n}^{a\rightarrow c}$, $\Delta\Psi_{n}^{a\rightarrow b}$, and $\Delta\Psi_{n}^{b\rightarrow c}$ represent the flow-plane twists, describing the variation of $\Psi_{n}$ between the regions $a\rightarrow c$, $a\rightarrow b$, and $b\rightarrow c$, respectively.
The observable defined in Eq. 2 is designed to measure the relative decrease in the anisotropic flow caused by the $\Psi_{n}$ decorrelation. This is evaluated over a “$2\eta$” distance (where $|\eta_{b}|=|\eta_{c}|=\eta$) between regions $b$ and $c$, with respect to a reference pseudorapidity window $a$. In practice, the decorrelation measured by Eq. 2 is independent of the reference pseudorapidity window $a$ only if there is no contribution from the term $\langle\Delta\Psi_{n}^{a\rightarrow b}\cdot\Delta\Psi_{n}^{b\rightarrow c}\rangle$ (see Eq. 3), i.e., if the consecutive changes in the flow-plane orientations $\Delta\Psi_{n}^{a\rightarrow b}$ and $\Delta\Psi_{n}^{b\rightarrow c}$ of the flow-plane angles are independent.
This paper presents an investigation of the longitudinal flow-plane decorrelations using a four-plane cumulant $T_{n}$ observable, proposed in Ref. Xu et al. (2022), using flow-plane angles from the backward, forward, and two mid-pseudorapidity regions. The $T_{n}$ observable is defined by
$T_{n}\{ba;dc\}=\langle\langle\,\sin[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\sin[n(\Psi_{n}^{d}-\Psi_{n}^{c})]\,\rangle\rangle,$ (5)
[…]
The $T_{n}$ observable is largely insensitive to correlations unrelated to the flow-plane orientation (i.e., non-flow correlations due to few-particle clusters such as jets, heavy resonance decays, local charge conservation, etc.), as it is based on the fourth-order cumulant Borghini et al. (2001); Voloshin et al. (2010); et al. (2002). For the same reason, it is also insensitive to the “random-walk” type of flow decorrelations. Global momentum conservation, although it does not contribute to harmonic correlations with $n>1$, arises from many different few-particle processes and is therefore also suppressed in higher-order cumulant calculations. Similarly, the ratio observable $r_{n}(\eta_{a\rightarrow c/b})$ in Eq. 2 is sensitive to the flow magnitude fluctuations. As the flow-plane fluctuations are smaller in events with the larger flow Bozek and Broniowski (2018), the measurements of $r_{n}(\eta_{a\rightarrow c/b})$ and $T_{n}$ are likely to be biased toward smaller values.
The flow-plane decorrelations are expected to be relatively small. In this case,
$T_{n}\{ba;dc\}\approx\langle\langle\,n\Delta\Psi^{a\rightarrow b}_{n}\cdot n\Delta\Psi^{c\rightarrow d}_{n}\,\rangle\rangle,$ (6)
where $\Delta\Psi^{a\rightarrow b}_{n}=\Psi_{n}^{b}-\Psi_{n}^{a}$ and $\Delta\Psi^{c\rightarrow d}_{n}=\Psi_{n}^{d}-\Psi_{n}^{c}$ are $\Psi_{n}$ twists from the pseudorapidity subevents $a\rightarrow b$ and $c\rightarrow d$. Therefore, $T_{n}$ represents the covariance between changes in orientation of flow plane from pseudorapidity window $a$ to $b$ and that from $c$ to $d$, quantifying the overall azimuthal twist of flow plane on moving forward in the pseudorapidity. One would expect a positive value of $T_{n}\{ba;dc\}$ when the flow-plane angles at backward and forward pseudorapidities are fluctuating in opposite directions to that at mid-pseudorapidity. On the other hand, the negative sign of $T_{n}\{ba;dc\}$ would correspond to the decorrelation pattern when the flow-plane angles at backward and forward pseudorapidities fall on the same direction relative to that at mid-pseudorapidity. These two distinct cases $T_{n}>0$ and $T_{n}<0$ are identified as the “S-shaped” and “C-shaped” decorrelations, respectively, in Ref. Xu et al. (2022).
In this analysis, we attempt to determine and identify the dominant pattern of the longitudinal flow-plane decorrelations by measuring $r_{n}$, $r_{n}^{\Psi}$, and $T_{n}$ observables using the flow-plane angles reconstructed in the forward, backward, and mid-pseudorapidity regions in isobar (Ru+Ru, Zr+Zr) and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV with the STAR detector at RHIC.
II EXPERIMENT AND DATA ANALYSIS
II.1 Datasets and Event Selection
Figure 1: Diagram illustrating four rapidity regions used for $T_{n}$ calculations.
The primary dataset used in this work was collected by the STAR experiment et al. (1994) for isobar Ru+Ru and Zr+Zr collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV in $2018$ to search for the Chiral Magnetic Effect (CME) et al. (2022). The main detectors used in the analysis are the Time Projection Chamber (TPC) et al. (STAR Collaboration), Event Plane Detector (EPD) et al. (STAR Collaboration), and Beam-Beam Counter (BBC) et al. (STAR Collaboration). To study the system size dependence of the $T_{n}$ observable, we also analyze the data of Au$+$Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV taken in years $2011$ and $2016$. The EPD subsystem used in the analysis of the isobar collisions is an detector upgrade installed in the STAR detector system in $2018$. It significantly improves the measurements of the event-plane angle in forward and backward pseudorapidity regions, compared to the BBC detector used earlier. Since the Au+Au collisions data in $2011$ and $2016$ were taken using the TPC and BBC, the BBC is used in the analyses of these datasets to estimate the event-plane orientations at forward and backward pseudorapidities. The results using Au+Au and isobar collisions are then corrected by corresponding BBC or EPD event plane resolutions.
Refer to caption Refer to caption
Figure 2: Left: The second harmonic event plane resolutions from the TPC, EPD, and BBC detectors. Right: The third harmonic sub-event plane resolutions from the TPC and EPD. The filled and open data points are for Ru$+$Ru and Zr$+$Zr collisions, respectively. The Zr$+$Zr data points are shifted towards right along the x-axis for clarity.
Many details of this analysis are similar to those in the CME isobar analysis. The event and the track selections, as described in details in Ref. et al. (2022), have been used on the same datasets for this analysis. The vertex distribution on isobar collisions exhibited asymmetry with a peak around $5$ cm due to online vertex selection. Therefore, an asymmetric vertex position cut of $-35<v_{z}<25$ cm on the primary $z-$vertex measured using a coordinate system with the origin at the TPC center was applied to maximize the statistics. Additional cut of $v_{r}=\sqrt{v_{x}^{2}+v_{y}^{2}}<2$ cm in the transverse plane was also applied for each collision event. In order to ensure the selection of good events, requirement of $|v_{z,\mathrm{TPC}}-v_{z,\mathrm{VPD}}|<5$ cm was imposed on the primary collision vertices, where $v_{z,\mathrm{VPD}}$ represents the primary interaction position along the beam direction measured by the Vertex Position Detectors (VPDs) et al. (2004). In addition, we require tracks reconstructed in the TPC to have a distance of closest approach (DCA) to the primary vertex to be less than $3$ cm. We also require each track to have greater than $15$ ionization points reconstructed in the TPC.
The symmetric vertex position cuts of $-30<v_{z}<30$ cm and $-6<v_{z}<6$ cm were applied to the $2011$ and $2016$ Au+Au collision datasets, respectively, as online vertex reconstruction for these collisions indicated symmetric $v_{z}$ distributions. Due to the deployment of the Heavy Flavor Tracker (HFT) Bouchet (2009) as an upgrade to the inner tracking system of the STAR experiment, a relatively narrow $v_{z}$ cut was applied for the $2016$ Au+Au collision dataset. Other track selection cuts used for the Au+Au collision analysis are same as those used for the isobar collision analysis.
II.2 Event Planes Reconstruction
The second- and third-order flow vectors $Q_{n}\equiv(Q_{n,x},Q_{n,y})$ and the event-plane angles $\Psi_{n}$ are reconstructed from the anisotropic flow itself following the procedure described in Refs. Poskanzer and Voloshin (1998); Voloshin et al. (2010):
$Q_{n,x}=\sum w_{i}\cos(n\phi_{i}),$ (7)
$Q_{n,y}=\sum w_{i}\sin(n\phi_{i}),$ (8)
$\Psi_{n}=\frac{1}{n}\arctan\left(\frac{Q_{n,y}}{Q_{n,x}}\right),$ (9)
where $w_{i}$ is the weight for $i^{th}$ particle, and the sum goes over all particles used in the calculation. For the flow vector and event-plane angle reconstruction in the TPC, primary tracks with $2.00.2<p_{\mathrm{T}}<$ 2.0 GeV/$c$ were taken from $-1<\eta<-0.1$ and $10.1<\eta<$ 1. The flow vectors were reconstructed with unit weight for the Particle-level calculations and with transverse momentum weight for the Q-level calculations (see below). Charged particles emitted in the forward and backward rapidity regions create a signal distribution with distinct peaks, each representing different numbers of minimally ionizing particles (nMIP) in the EPD, BBC tiles. In the EPD, the event-plane angles were reconstructed using nMIP of each EPD tile $i$ which satisfies $0.3\leq$ nMIP $\leq 3$ as weight, $w_{i}$, in Eqs. 7 and  8. The weight is set to $3$ if nMIP $>3$. Similar information is used to reconstruct event-plane angles in the BBC.
Recentering of flow vectors and flattening of $\Psi_{n}$ distribution were performed to remove any acceptance effects. This procedure was done on the run-by-run basis in each z-vertex bin of $2$ cm width and centrality bin of $5\%$ step.
Figure 2 shows the event-plane resolutions as a function of centrality for the second- and third-order anisotropic flow calculated for the TPC, EPD, and BBC pseudorapidity subevents in the isobar collisions. The event-plane resolution calculations were done with the two- and three-subevent methods Voloshin et al. (2010).
II.3 Calculation of the Four-Plane Correlator
Refer to captionRefer to caption
Figure 3: (Color online) Resolution ${\rm Res}(T_{n}\{ba;dc\})$ plots for the Q-level (left) and Particle-level (right) calculations using Ru$+$Ru (solid markers) and Zr$+$Zr (open markers) collisions for the second-order anisotropic flow. The Zr$+$Zr data points are shifted towards right along x-axis for clarity.
In this work, the measurements of the $T_{n}\{ba;dc\}$ correlators for the second- and third-order anisotropic flow are performed using flow-plane angles reconstructed in the TPC and EPD (or BBC) subsytems of the STAR detector. Therefore, in the measurements presented in this paper, the pseudorapidity windows $a$, $b$, $c$, and $d$ from the backward, mid-, and forward pseudorapidity regions as depicted in Fig. 1 correspond to the EPD-east ($-5.1<\eta<-2.1$) or BBC-east ($-5.0<\eta<-3.3$), TPC-east ($-1.0<\eta<-0.1$), TPC-west ($1.00.1<\eta<$ 1.0), and EPD-west ($5.12.1<\eta<$ 5.1) or BBC-west ($5.03.3<\eta<$ 5.0) regions, respectively.
Refer to captionRefer to caption
Figure 4: (Color online) Same as in Fig. 3 for the third-order anisotropic flow.
The measurement using four flow planes from the east and west regions of the TPC and EPD (or BBC) is referred as the Q-level calculation, which follows Eq. 10. Similarly, the measurement using two flow planes from the east and west regions of the EPD (or BBC) along with the particle track’s azimuth $\phi$ from the east and west regions of the TPC is identified as the Particle-level calculation. The Particle-level calculation follows Eq. 11 for the $T_{n}\{ba;dc\}$ observable.
Experimentally, the four-plane correlator observable, $T_{n}$, can be calculated as
[…]
for the Q-level calculation and
[…]
for the Particle-level calculation.
The observed cumulant $T^{obs}_{n}\{ba;dc\}$ in the numerator of Eq. 10 is calculated using four $n^{th}$-harmonic event planes from the backward, mid-, and forward pseudorapidity regions $a$, $b$, $c$, and $d$ as shown in Fig. 1, which has to be corrected for the event-plane resolution. The term ${\rm Res}(T_{n}\{ba;dc\})$ in the denominator is used to correct $T^{obs}_{n}\{ba;dc\}$ measurement for the event-plane resolution.
Refer to captionRefer to caption
Figure 5: (Color online) Centrality dependence of different terms contributing to $T^{obs}_{2}\{ba;dc\}$, as given in Eqs. 19 and 20: Solid markers show Ru$+$Ru and open markers Zr$+$Zr results for the Q-level (left panel) and Particle-level (right panel) calculations. The Zr$+$Zr data points are shifted towards right for clarity.
Refer to captionRefer to caption
Figure 6: Same as Fig. 5 for the third-order anisotropic flow $n=3$.
Since Eq. 10 utilizes the event planes reconstructed from the four different pseudorapidity regions, ${\rm Res}(T_{n}\{ba;dc\})$ in general can be expressed as the product of the four event-plane resolutions,
${\rm Res}(T_{n}\{ba;dc\})={\rm Res}(\Psi^{a}_{n}){\rm Res}(\Psi^{b}_{n}){\rm Res}(\Psi^{c}_{n}){\rm Res}(\Psi^{d}_{n}).$ (12)
In practice, the analysis estimates the ${\rm Res}(T_{n}\{ba;dc\})$ term using formula
$\begin{split}{\rm Res}(&T_{n}\{ba;dc\})\\ &=\langle\cos[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\rangle\langle\cos[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle,\end{split}$ (13)
for the Q-level calculations by identifying the following correlations
$\langle\cos[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\rangle$ $={\rm Res}(\Psi^{a}_{n}){\rm Res}(\Psi^{b}_{n}),$ (14)
$\langle\cos[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle$ $={\rm Res}(\Psi^{c}_{n}){\rm Res}(\Psi^{d}_{n}),$ (15)
as the product of the two event-plane resolutions. Using Eqs. 14 and 15, the product of the four event-plane resolutions in Eq. 12 follows Eq. 13. In the Particle-level calculations, the event-plane resolution correction ${\rm Res}(T^{p-level}_{n}\{ba;dc\})$ is estimated using azimuthal angles ($\phi^{b}$, $\phi^{c}$) of tracks in mid-pseudorapidity windows $b$ and $c$ instead of $\Psi^{b}_{n}$ and $\Psi^{c}_{n}$ in Eq. 13, i.e., using formula
$\begin{split}{\rm Res}(&T^{p-level}_{n}\{ba;dc\})\\ &=\langle\cos[n(\phi^{b}-\Psi^{a}_{n})]\rangle\langle\cos[n(\Psi_{n}^{d}-\phi^{c})]\rangle\\ &=v_{n}(\eta_{b})v_{n}(\eta_{c}){\rm Res}(\Psi^{a}_{n}){\rm Res}(\Psi^{d}_{n}),\end{split}$ (16)
where $v_{n}(\eta_{b})$ and $v_{n}(\eta_{c})$ represent the $n^{th}$-order flow coefficients such that
$\langle\cos[n(\phi^{b}-\Psi^{a}_{n})]\rangle$ $=v_{n}(\eta_{b}){\rm Res}(\Psi^{a}_{n}),$ (17)
$\langle\cos[n(\phi^{c}-\Psi^{d}_{n})]\rangle$ $=v_{n}(\eta_{c}){\rm Res}(\Psi^{d}_{n}).$ (18)
The set of plots in Fig. 3 shows the centrality dependence of ${\rm Res}(T_{n}\{ba;dc\})$ and ${\rm Res}(T^{p-level}_{n}\{ba;dc\})$ for the elliptic flow in isobar collisions. A similar set of plots for triangular flow in isobar collisions are shown in Fig. 4.
In the following, the centrality dependence of the observed four-plane cumulant $T^{obs}_{n}\{ba;dc\}$ and its contributing components are discussed. The $T^{obs}_{n}\{ba;dc\}$ correlator, explicitly derived in Ref. Xu et al. (2022), consists of three terms:
$\begin{split}&T^{obs}_{n}\{ba;dc\}\\ &=\langle\langle\sin[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle\rangle\\ &\equiv\langle\sin[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle\\ &-\frac{1}{2}\langle\cos[n(\Psi^{b}_{n}-\Psi^{d}_{n})]\rangle\langle\cos[n(\Psi^{a}_{n}-\Psi^{c}_{n})]\rangle\\ &+\frac{1}{2}\langle\cos[n(\Psi^{a}_{n}-\Psi^{d}_{n})]\rangle\langle\cos[n(\Psi^{b}_{n}-\Psi^{c}_{n})]\rangle,\end{split}$ (19)
[…]
$\begin{split}&(T^{obs}_{n}\{ba;dc\})^{p-level}\\ &=\langle\langle\sin[n(\phi^{b}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\phi^{c})]\rangle\rangle\\ &\equiv\langle\sin[n(\phi^{b}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\phi^{c})]\rangle\\ &-\frac{1}{2}\langle\cos[n(\phi^{b}-\Psi^{d}_{n})]\rangle\langle\cos[n(\Psi^{a}_{n}-\phi^{c})]\rangle\\ &+\frac{1}{2}\langle\cos[n(\Psi^{a}_{n}-\Psi^{d}_{n})]\rangle\langle\cos[n(\phi^{b}-\phi^{c})]\rangle,\end{split}$ (20)
Figures 5 and 6 show the centrality dependence of different terms contributing to the $T^{obs}_{n}\{ba;dc\}$ observable for isobar collisions. Results from the Q-level and Particle-level calculations (as discussed in Sec. II.3) are presented side-by-side for better comparison. The centrality dependence of different terms from the Q-level and Particle-level approaches is slightly different. This effect is similar to the one responsible for the difference in the centrality dependence of the event-plane resolution and elliptic flow.
Refer to captionRefer to caption
Figure 7: Ratio observable $r_{n}$ (defined in Eq. 22) as a function of centrality for the second- (left) and third-order (right) anisotropic flow. (Color online) Red circle and blue star data points correspond, respectively, to Ru$+$Ru and Zr$+$Zr collisions. The systematic uncertainties are indicated by boxes. The Zr+Zr data points are shifted towards right for clarity.
II.4 Systematic Uncertainty
In this section, we discuss different sources of systematic uncertainty affecting the $T_{n}$ and $r_{n}$ measurements and the methodology implemented to estimate the total systematic uncertainty. Variations in analysis cuts such as those on the $v_{z}$ selection, DCA, number of reconstructed TPC hits (nHitsFit), and the TPC $\eta$-gap are considered major sources contributing to the total systematic uncertainty in our measurements. All individual contributions to the total systematic uncertainty have been calculated in each centrality bin of $5\%$ step. A summary of the nominal and altered cut selections that are to estimate the systematic uncertainty in this analysis is provided below.
[…]
  • Altered cut selection 1: $-35<v_{z}<0$ cm (all other cuts are the same as in the nominal setup).
  • Altered cut selection 2: DCA $<2$ cm (all other cuts are the same as in the nominal setup).
  • Altered cut selection 3: nHitsFit $>20$ (all other cuts are the same as in the nominal setup).
[…]
The overall systematic uncertainty is then calculated by adding the individual contributions from different sources in quadrature assuming that the individual systematic uncertainties are uncorrelated. The Barlow approach Barlow (2002) that was used in the CME search analysis is implemented to avoid the statistical fluctuations. The overall procedure is summarized below:
  1. 1.
    Calculate the difference in the observable $Y_{n}$ (where $Y_{n}$ represents either $r_{n}$, $r_{n}^{\Psi}$, or $T_{n}$; see Secs. II.3 and III for definitions) when computed using the nominal cut and each of the altered cuts: $\Delta Y^{i}=Y^{\text{alt},i}_{n}-Y^{\text{nom}}_{n}$.
  2. 2.
    Calculate the effect of statistical fluctuation defined as $\Delta E_{i}=\sqrt{|E_{\text{alt},i}^{2}-E_{\text{nom}}^{2}|}$, where $E$ is the statistical uncertainty of $Y_{n}$ measurement.
  3. 3.
    Calculate the systematic uncertainty contribution from each source as $\sigma_{i}=\sqrt{{\Delta Y^{i}}^{2}-{\Delta E_{i}}^{2}}$, but set $\sigma_{i}=0$ if $|\Delta Y^{i}|<\Delta E_{i}$, i.e., if the change $|\Delta Y^{i}|$ is consistent with statistical fluctuations.
  4. 4.
    Estimate the total systematic uncertainty by adding the individual contributions from different sources in quadrature using formula
    $\sigma_{\text{tot}}=\sqrt{\sum^{N}_{i=1}\sigma_{i}^{2}},$
    where $N=4$, corresponding to four measurements from different cut variations.
Finally, if the systematic uncertainty estimates obtained using the procedure described above fluctuate across centrality bins, a polynomial fit is applied to smooth the total systematic uncertainty as a function of centrality.
Refer to captionRefer to caption
Figure 8: Comparison of ratio observables $r_{n}$ and $r_{n}^{\Psi}$ (defined in Eqs. 22 and 23) as a function of centrality for the second (left) and third (right) harmonic anisotropic flow in Ru$+$Ru and Zr$+$Zr collisions. The systematic uncertainties are indicated by boxes. (Color online) Red circles correspond to the Particle-level calculation, and blue stars (shifted horizontally for clarity) correspond to the Q-level calculations (see text for details).
III Results
III.1 Ratio Observables
Calculation of the ratio observable involves measuring anisotropic flow over different regions of interest with respect to a reference window. We calculated the observable $r_{n}(\eta_{a\rightarrow c/b})$ defined in Eq. 2 for $n=2$ and $3$ by measuring the anisotropic flow in the TPC-west to that in the TPC-east with respect to the fixed reference rapidity window EPD-east.
The similar ratio $r_{n}(\eta_{d\rightarrow b/c})$ with the EPD-west as a reference rapidity window is calculated using,
$r_{n}(\eta_{d\rightarrow b/c})=\frac{\langle\cos[n(\phi^{b}-\Psi^{d}_{n})]\rangle}{\langle\cos[n(\phi^{c}-\Psi^{d}_{n})]\rangle}.$ (21)
Taking advantage of the symmetric STAR detector system, we combined these two measurements, $r_{n}(\eta_{a\rightarrow c/b})$ and $r_{n}(\eta_{d\rightarrow b/c})$, to improve statistical precision in the final result using,
$r_{n}=\frac{1}{2}[r_{n}(\eta_{a\rightarrow c/b})+r_{n}(\eta_{d\rightarrow b/c})].$ (22)
Additionally, the decorrelations of the flow-plane angle are also calculated using the event-plane angles $\Psi_{n}$ in the ratio formulation using,
$\begin{split}r^{\Psi}_{n}&=\frac{1}{2}[r^{\Psi}_{n}(\eta_{a\rightarrow c/b})+r^{\Psi}_{n}(\eta_{d\rightarrow b/c})]\\ &=\frac{1}{2}\bigg{[}\frac{\langle\cos[n(\Psi^{c}_{n}-\Psi^{a}_{n})]\rangle}{\langle\cos[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\rangle}+\frac{\langle\cos[n(\Psi^{b}_{n}-\Psi^{d}_{n})]\rangle}{\langle\cos[n(\Psi^{c}_{n}-\Psi^{d}_{n})]\rangle}\bigg{]}.\end{split}$ (23)
The measurements of $r_{n}$ and $r^{\Psi}_{n}$ correspond to the factorization ratios obtained through the Particle-level and Q-level methods discussed earlier. Note that the Q-level calculation utilizes $\Psi_{n}$’s from the TPC, reconstructed using $p_{\mathrm{T}}$ weights, resulting in better event-plane resolutions.
Figure 7 presents the results for the factorization ratio $r_{n}$ as a function of centrality for the second and third harmonic flow in isobar collisions. The $r_{3}$ results for $60-80\%$ centrality range have large uncertainties due to low multiplicity and are not shown. In mid-peripheral collisions, small differences in the $r_{n}$ measurements are observed between the Ru+Ru and Zr+Zr collision systems.
Refer to captionRefer to caption
Figure 9: (Color online) $T_{n}\{ba;dc\}$ observable as a function of centrality for the elliptic (left) and triangular (right) flow in Ru+Ru and Zr+Zr collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. The systematic uncertainties are indicated by boxes. Red circles correspond to the Particle-level calculation, and blue stars (shifted horizontally for clarity) correspond to the Q-level calculations.
The ratios $r_{n}$ and $r^{\Psi}_{n}$ are less than unity as expected due to decorrelations increasing further away from the reference rapidity window. Both the $r_{n}$ and $r^{\Psi}_{n}$ results are found to be in good agreement with each other within experimental uncertainties at all centralities despite our use of flow vectors reconstructed using different weights in the two measurements. The comparison plot of $r_{n}$ and $r^{\Psi}_{n}$ is presented for the combined isobar data rather than for the individual Ru+Ru and Zr+Zr data sets, as shown in Fig. 8.
The centrality dependence of $r_{n}$ qualitatively can be explained as follows. In the mid-central collision, the flow is large and the effect of fluctuations is relatively small, resulting in the $r_{2}$ values closer to unity. On the other hand, in central collisions the flow is small and thus the relative effect of fluctuations increases. In peripheral (low multiplicity) collisions the flow is decreasing while the fluctuations become larger due to the small number of participants. The $r_{3}$ is the largest in most central collisions where the triangular flow becomes the strongest. Similar reasoning applies to explain the feature of $r^{\Psi}_{n}$ as a function of centrality.
Figure 8 shows comparison between $r^{\Psi}_{n}$ (Eq. 23, Q-level) and $r_{n}$ (Eq. 22, Particle-level) measurements for the second- and third-order flow in isobar collisions. The combined dataset of Ru+Ru and Zr+Zr collisions is used to improve the statistical precision. Using $\Psi_{n}$’s with improved resolutions due to $p_{\mathrm{T}}$ weighting, one might intuitively anticipate relatively smaller decorrelation in the Q-level measurements. However, Fig. 8 shows a negligible difference between the Q-level and Particle-level ratio measurements in each centrality bin, despite the use of $p_{\mathrm{T}}$ weights in the former. This clearly indicates that our ratio measurements primarily reflect the dynamic changes rather than statistical fluctuations. These dynamic changes encompass contributions from both the random-walk and the long-range systematic variations of the flow plane, as discussed in Sec. I. Such variations are believed to be stem out from the initial state fluctuations of the collision systems.
Figure 10: Same as Fig. 9 but using forward and backward flow planes from BBCs instead of EPD for isobar collisions and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. Green diamond data points correspond to the Q-level calculations in Au+Au collisions.
III.2 $T_{n}$ Observables
Previous subsections reported the measurements of observed $T^{obs}_{n}\{ba;dc\}$ and corresponding event-plane resolution correction terms for the elliptic and triangular flow for isobar collisions. The final results for the $T_{n}$ observable, corrected for the event-plane resolution, are discussed in this section. Regardless of the small difference Xu et al. (2018) in the nuclear deformation of the two isobar species, the results shown in previous sections are consistent within the statistical uncertainties for Ru+Ru and Zr+Zr collisions. Hence, the final results are calculated using the combined Ru+Ru and Zr+Zr collision data for a better statistical precision over the wider range of centrality.
Figures 9 and 10 present the results of $T_{n}\{ba;dc\}$ measurements as a function of centrality for the elliptic and triangular flow in isobar collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. The results for the $60-80\%$ centrality range have large uncertainties due to low multiplicities and are not shown. The comparison between the results obtained using the Q-level and Particle-level approaches has been shown for a centrality range of $0-60\%$ in $12$ centrality bins. The results from the two approaches are in good agreement with each other within experimental uncertainties. The $T_{2}$ (Q-level) measurement averaged over the $10-30\%$ centrality range is $-0.004\pm 0.001(\rm stat)\pm 0.002(\rm syst)$ and consistent with zero, within experimental uncertainties, for other centralities. The $T_{3}$ result is consistent with zero within experimental uncertainties in all centrality bins. Such small values of $T_{n}$ observables from experimental data support a “random-walk” variation of flow-plane angles with rapidity (see Sec. IV for detailed discussion and summary).
The similar measurements for the elliptic flow in Au+Au and isobar collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV using four flow planes reconstructed from the positive and negative pseudorapidity sides of the TPC, and BBC are shown in Fig. 10. Within the statistical uncertainty, the $T_{n}$ measurements in Au+Au and isobar collisions are found to be compatible with each other indicating weak dependence of the results on the system size.
Our $T_{n}$ measurements are in agreement, within the statistical and systematic uncertainties, with the AMPT model calculations from Ref. Xu et al. (2022) for Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. The study claims that the $T_{n}$ observable is largely insensitive to various non-flow effects and can effectively distinguish between different decorrelation patterns. The calculation reported in Ref. Xu et al. (2022) yields slightly positive $T_{2}$ values, approximately on the order of $10^{-3}$ for events within the $20-40\%$ centrality range. Due to differences in the ordering of the flow-plane angles, a positive $T_{2}$ as defined in Ref. Xu et al. (2022) corresponds to a negative $T_{2}$ in our convention.
IV Discussions and summary
We measured the four-plane cumulant correlator $T_{n}$, and the factorization ratios $r_{n}$ and $r_{n}^{\Psi}$, in isobar (Ru+Ru and Zr+Zr) and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. These measurements focus on $n=2$ and $n=3$ flow harmonics, testing the longitudinal flow-plane decorrelation phenomena in heavy-ion collisions.
In mid-central collisions ($10-30\%$ centrality range), $T_{2}\{ba;dc\}=-0.004\pm 0.001(\mathrm{stat})\pm 0.002(\mathrm{syst})$ have been observed independent of the collision system and consistent with zero within experimental uncertainties at other centralities. The $T_{3}\{ba;dc\}$ result is consistent with zero within experimental uncertainties at all centrality bins. The factorization ratio observables $r^{\Psi}_{n}$ and $r_{n}$ are found to be similar within experimental uncertainties across all centralities, despite the use of $p_{T}$ weights in the former. The $r_{2}$ observable for $n=2$ flow harmonics reaches its maximum value of $r_{2}\approx 0.96$ in the $10-30\%$ centrality range. Similarly, the maximum value of $r_{3}\approx 0.95$ is observed in the most central $0-5\%$ centrality bin. As outlined in Sec. I, such values of the factorization ratio correspond to flow decorrelations of approximately $\frac{n^{2}}{2}\langle(\Delta\Psi^{b\rightarrow c}_{n})^{2}\rangle\sim 0.04$ and $\sim 0.05$ for $n=2$ and $n=3$ harmonics, respectively. Considering these decorrelation estimates, if one assumes a similar order of decorrelation between mid- and forward (backward) pseudorapidities, the prediction for the $T_{n}\{ba;dc\}\sim n^{2}\langle\Delta\Psi^{a\rightarrow b}_{n}\cdot\Delta\Psi^{c\rightarrow d}_{n}\rangle$ observable would be at least $10$ times larger in magnitude than those observed.
Additionally, the explicitly measured values of $T_{n}$ in this paper are significantly smaller than what one would expect in the scenario of monotonic variation of the flow angles between forward and backward rapidity windows determined by eccentricities calculated using participating projectile and target nuclei. According to calculations in Ref. Jia and Huo (2014b), the RMS value of the forward-backward twist $(2\Psi_{B}-2\Psi_{F})$ distribution in such an approach is about $0.5$ for mid-central collisions, which would translate to $T_{2}\approx 0.5^{2}/4$, again much larger than values we observed experimentally.
As noticed in Sec. I, flow magnitude fluctuations likely bias the measurements of $r_{n}$ and $T_{n}$ correlators toward smaller values, but this effect would not be able to explain an order of magnitude difference in the above estimates. In the future high statistics measurements, it will be possible to study the effect of flow magnitude fluctuations by introducing additional $v_{n}$ weights in the definitions of Eq. 5.
The above considerations lead to a conclusion that our results favor a picture of a “random-walk” variation of the flow-plane angles, with a rapidity correlation length smaller than the rapidity region under study. In this scenario, the variation of the flow-plane angle along rapidity is dominated by random fluctuations with the correlation length in rapidity being much smaller than the size of rapidity windows used in this analysis. These type of decorrelations contribute to the ratio observable $r_{n}$ but not to $T_{n}$ due to the cumulant character of the latter. As was noted in Ref. Xu et al. (2022), this type of decorrelation would also lead to almost no dependence of $r_{n}$ on the reference rapidity, which was observed experimentally. Slightly negative result for $T_{n}$ could indicate that both flow planes at mid-pseudorapidity are fluctuating together relative to those at the forward and backward pseudorapidities. These results provide new insights into the decorrelation patterns within the system and offer a quantitative estimate of possible systematic variations in anisotropic flow angles - such as the “twist” between forward and backward pseudorapidity regions. The study opens new avenues for exploring the three-dimensional structure of the QGP formed in heavy-ion collisions.
Acknowledgements.
We thank the RHIC Operations Group and SDCC at BNL, the NERSC Center at LBNL, and the Open Science Grid consortium for providing resources and support. This work was supported in part by the Office of Nuclear Physics within the U.S. DOE Office of Science, the U.S. National Science Foundation, National Natural Science Foundation of China, Chinese Academy of Science, the Ministry of Science and Technology of China and the Chinese Ministry of Education, NSTC Taipei, the National Research Foundation of Korea, Czech Science Foundation and Ministry of Education, Youth and Sports of the Czech Republic, Hungarian National Research, Development and Innovation Office, New National Excellency Programme of the Hungarian Ministry of Human Capacities, Department of Atomic Energy and Department of Science and Technology of the Government of India, the National Science Centre and WUT ID-UB of Poland, the Ministry of Science, Education and Sports of the Republic of Croatia, German Bundesministerium für Bildung, Wissenschaft, Forschung and Technologie (BMBF), Helmholtz Association, Ministry of Education, Culture, Sports, Science, and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) and Agencia Nacional de Investigación y Desarrollo (ANID) of Chile.